An effective rate of 29.61% corresponds to a nominal rate,compounded weekly,of?

1
Expert's answer
2022-03-04T05:38:22-0500

Solution


The effective annual interest rate is calculated using the following formula: "j = {\left( {1 + \frac{i}{n}} \right)^n} - 1"


Here

j - is the effective annual interest rate,

i - is the nominal interest rate,

n - is the number of periods.


Given that effective annual interest rate "j = 29.61%" %

And when compounded weekly, "n=52" (52 weeks in a year)

 

Solving "j = {\left( {1 + \frac{i}{n}} \right)^n} - 1" for "i"


"j = (1+\frac{i}{n})^{n}-1\\"


"j+1=(1+\frac{i}{n})^{n}\\"


"(j+1)^{\frac{1}{n}} = (1+\frac{i}{n})\\"


"\frac{i}{n}=(j+1)^{\frac{1}{n}}-1\\"


"i=n\times \left [ (j+1)^{\frac{1}{n}}-1 \right ]"




Replacing "j = 29.61%" % and "n=52"


"i=52\times \left [ (\frac{29.61}{100}+1)^{\frac{1}{52}}-1 \right ]"


"i=0.2600076349\"


"i=26" %



Hence the nominal rate will be "i=26" %